# Cross-lingual Contextualized Phrase Retrieval

### v1

@@ Abstract
Phrase-level dense retrieval has shown many appealing characteristics in downstream NLP tasks by leveraging the fine-grained information that phrases offer. In our work, we propose a new task formulation of dense retrieval, cross-lingual contextualized phrase retrieval, which aims to augment cross-lingual applications by addressing polysemy using context information. However, the lack of specific training data and models are the primary challenges to achieve our goal. As a result, we extract pairs of cross-lingual phrases using word alignment information automatically induced from parallel sentences. Subsequently, we train our Cross-lingual Contextualized Phrase Retriever (CCPR) using contrastive learning, which encourages the hidden representations of phrases with similar contexts and semantics to align closely. Comprehensive experiments on both the cross-lingual phrase retrieval task and a downstream task, i.e, machine translation, demonstrate the effectiveness of CCPR. On the phrase retrieval task, CCPR surpasses baselines by a significant margin, achieving a top-1 accuracy that is at least 13 points higher. When utilizing CCPR to augment the large-language-model-based translator, it achieves average gains of 0.7 and 1.5 in BERTScore for translations from XEn and vice versa, respectively, on WMT16 dataset. Our code and data are available at https://github.com/ghrua/ccpr_release.
@@ 1 Introduction
Therefore, we advocate for a new task formulation, i.e., cross-lingual contextualized phrase retrieval, which aims to find the cross-lingual phrase that is mostly relevant to the provided (general-type) source phrases, considering their meanings and surrounding contexts. However, achieving this goal is non-trivial, due to the scarcity of specific training data and models. Since annotating cross-lingual contextualized phrase pairs of general type is very difficult and expensive, we first introduce a data collection method, which leverages the automatically induced word alignment information from parallel sentences to extract suitable cross-lingual phrase pairs for training. This ensures that the phrases are of general type and both cross-lingual phrases and contexts are well aligned. Thereafter, we propose a Cross-lingual Contextualized Phrase Retriever (CCPR) which is trained on the constructed dataset. The CCPR mainly employs phrase-level contrastive learning to draw cross-lingual phrases with similar contexts and meanings closer in the hidden space. At inference time, we can leverage the well-trained CCPR to build a phrase-level index and use the query phases for search. One remaining question is how to select meaningful phrases for indexing. To address this problem, we propose to learn a phrase segmentation module, which can be used to predict meaningful phrases from sentence- or passage-level retrieval data at inference time. Both of the modules are critical for employing our CCPR in real-word applications.
@@
Experiments on both of the cross-lingual contextualized phrase retrieval and a downstream task, i.e., machine translation, show the effectiveness of CCPR. We first evaluate CCPR on the cross-lingual contextualized phrase retrieval task, since a higher accuracy is more beneficial for downstream tasks, such as machine translation, cross-lingual dictionary induction, etc Zhang and Zong (2016); Søgaard et al. (2018). For this task, we use the human annotated cross-lingual phrase pairs as the golden truth and evaluate whether CCPR and other baselines can retrieve those golden phrases from a large-scale index, which contains approximately 10M phrases. In the experiments, CCPR outperforms the baselines for at least 13 points in terms of the top-1 accuracy. We also conduct evaluation on the machine translation task. The purpose of machine translation task is to evaluate whether the information retrieved by CCPR is beneficial for cross-lingual generation. As a result, we simply integrate the retrieved phrase information to the input of an large language model (LLM), e.g., LLama-2 Touvron et al. (2023), and compare our method with other baselines. For both XEn and EnX translation directions on the WMT16 dataset, where X is from six languages, our CCPR achieves averaged gians of 0.7 and 1.5 BERTScore points, respectively.
@@
We propose a Cross-lingual Contextualized Phrase Retriever (CCPR), which leverages the constructed training data, i.e., cross-lingual phrase pairs extracted from automatically induced word alignment information, to learn both the phrase alignment and segmentation modules.
We conduct extensive experiments on cross-lingual contextualized phrase retrieval and machine translation. Our method outperforms the baselines by a large margin.
@@ 2 Related Work
In the literature of text generation, dense retrieval has been widely studied at both sentence and phrase levels. Particularly after the emergence of large language models (LLMs), how to use dense retrieval to retrieve related sentences (or passages) Izacard et al. (2021); Ni et al. (2022) for input to LLMs has became a popular research topic, namely, retrieval-augmented generation (RAG) Lewis et al. (2020b); Karpukhin et al. (2020); Guu et al. (2020); Borgeaud et al. (2022); Asai et al. (2024). However, compared to the sentence-level retrieval, phrase-level dense retrieval has been shown to not only enhance retrieval accuracy but also markedly boost performance in downstream tasks like open-domain question answering (QA) Lee et al. (2021a, b). By indexing and integrating retrieved phrases into the output of language models, Lan et al. (2023) and Cao et al. (2024) make the text generation process more attributable and accurate. These studies underline the effectiveness of phrase-level dense retrieval. However, our work notably diverges from those research works, because they focus more on the English-only tasks.
@@
In cross-lingual tasks, the majority of research has concentrated on sentence-level retrieval. Many research works in retrieval-augmented machine translation Zhang et al. (2018); Gu et al. (2018); Xia et al. (2019); He et al. (2021) have explored retrieving bilingual sentences based on similarity of source sentences, guiding translation models to utilize retrieved target information. A notable limitation of existing approaches within this research domain is their reliance on bilingual data for constructing the retrieval index. This requirement substantially limits the scale and diversity of data available for retrieval purposes. In contrast, Cai et al. (2021) propose to directly retrieving relevant target sentences using source sentences to assist machine translation, facilitating cross-lingual retrieval using a monolingual index. Different from the focus of Cai et al. (2021), which aims for a task-specific model for machine translation, numerous studies in dense retrieval aim to enhance the cross-lingual sentence retrieval on more languages and at a larger scale Chidambaram et al. (2019); Conneau et al. (2020); Reimers and Gurevych (2020); Heffernan et al. (2022); Feng et al. (2022); Cai et al. (2022).
@@
Additionally, there is research focused on employing nearest neighbor retrieval to support machine translation Khandelwal et al. (2021); Zheng et al. (2021); Meng et al. (2022); Deguchi et al. (2023). These efforts differ from the aforementioned studies as they aim to estimate prediction distributions through nearest neighbor retrieval at each translation step.
Mining of Cross-lingual Phrase Pairs
The extraction of cross-lingual phrase pairs by utilizing the word alignment information has been a standard of the Phrase-based Statistical Machine Translation (PBSMT) Koehn et al. (2003); Chiang (2005). However, the phrase table extracted in PBSMT is static and context-independent, where each source phrase is mapped to multiple target phrases without considering the contexts. Similarly, Zheng et al. (2022) extract cross-lingual phrase pairs based on link information of wiki entities, which overlooks the nuanced meanings that phrases may convey in different contexts. Given the long-recognized importance of context in resolving polysemy within NLP Devlin et al. (2019), our training data and model are crafted to be context-dependent, offering greater utility for downstream NLP applications.
@@ 3 Task Formulation
Our research aims a cross-lingual phrase retrieval system that can augment downstream NLP tasks. To ensure its effectiveness, the retrieval system must adhere to two essential criteria. First, it should ensure that phrases in the retrieval index are of general type. Second, it must resolve the polysemy of general-type phrases by considering their contextual differences.
@@
This new task formulation also introduces several challenges, with the primary issue being the lack of training data. Specifically, there is a scarcity of cross-lingual phrase pairs of general type and accompanied by contextual information. Such data are crucial for training the model to recognize cross-lingual phrases that have similar contexts and meanings, allowing it to align these phrases closely in a hidden space.
@@ 4 Training Data Collection
Formally, given a pair of parallel sentences , and , where and are the number of words in the sequences, we use a word alignment model to obtain the word alignment information of and Koehn et al. (2003); Och et al. (1999). If every word in a consecutive span can be aligned to a consecutive span , we will use them as a pair of aligned phrases and , where and . Notably, phrases that meet the constraints may range from a single word to the whole sequence, and extracted phrases are allowed to have overlaps. In addition, the context information of and are and , respectively.
As shown in Figure 1 (a), The Germany phrase (“einer”, “Schweigeminute”) is aligned with the the English phrase (“a”, “minute”, “’s”, “silence”) within similar contexts, i.e., parallel sentences. More implementation details about the extraction of cross-lingual contextualized phrase pairs are discussed in Section 6.1.
@@ 5.1 Model Architecture
We introduce a Cross-lingual Contextualized Phrase Retriever (CCPR) for our new formulated task. The main target of CCPR is to make the representations of cross-lingual phrases with similar contexts and meanings to be close in the hidden space. To this end, we propose a cross-lingual phrase alignment module based on contrastive learning, utilizing the data collected in Section 4 for training. In addition, one remaining problem is how to select phrases for indexing at inference time. To address this, we propose a phrase segmentation module that uses phrase representations to predict meaningful phrases from sentences or paragraphs. At inference time, we can leverage the learned phrase segmentation module to select phrases for indexing, ensuring the train-test consistency.
@@
where is a Transformer model, e.g., BERT Devlin et al. (2019); Feng et al. (2022) and RoBERTa Conneau et al. (2020), that encodes the context to a matrix with a dropout mask . The and are the hidden representations of word and , respectively, i.e., the start and end token of the phrase. For example, in the phrase (“a”, “minute”, “’s”, “silence”), and . is a function that maps the concatenation of two hidden states and from to , where is the output hidden size. The encoding for phrases in context is similar.
@@
For a batch of parallel sentences, we can extract a collection of cross-lingual phrase pairs , where and are the positive examples of each other. Since each positive phrase pair is from parallel sentences that share similar semantics, directly learning on them may cause the model to learn a trivial shortcut. Therefore, inspired by SimCSE Gao et al. (2021), we apply two independently sampled dropout mask and , to encode phrases and , respectively, following Equation 1. The training objective for the direction is:
where we use all the non-paired phrases in as in-batch negatives in Equation 3. For bidirectional symmetry, our final loss is:
The illustration of the contrastive learning is shown in Figure 1 (a). In practice, we may use parallel data from multiple languages during training, i.e., the source contexts in a batch may come from multiple languages, e.g., Germany, Czech, etc, which is inspired by the success of multi-lingual training in Liu et al. (2020).
@@
This module aims to learn how to select phrases from sentences or passages, since most retrieval data at inference time comes in sentences or passages, rather than phrases. To ensure the train-test consistency, we take the phrases extracted in Section 4 as positive data, and all the other spans in the corresponding contexts as negative data. Our phrase segmentation module is a binary classifier defined as follows:
where is the label for the span , where , is an activation function, is linear layer that maps the hidden representation from to , and the definitions for and are the same as in Equation 1. We use the label for phrases and for non-phrase spans. The phrase segmentation for sequence is similar to Equation 5. In practice, the number of non-phrase spans is significantly more than the number of phrases. To mitigate data imbalance, we employ a strategy of randomly sampling an equal number of non-phrase spans and phrases within a sentence during training Li et al. (2020). The illustration of the phrase segmentation is shown in Figure 1 (b).
@@ 5.2 Inference Pipeline
The pipeline of employing our cross-lingual contextualized phrase retriever for downstream NLP tasks involves two main steps: index building and searching. An illustration of our inference pipeline is shown in Figure 1 (c).
@@
A significant benefit of our retriever is its ability to leverage vast monolingual data in the target language, eliminating the need for bilingual data to construct the index at inference time Zhang et al. (2018); Gu et al. (2018); Bapna and Firat (2019). Given the retrieval data in target language, our retriever begins by segmenting sentences (or passages) into phrases. When segmenting a sentence (or passage), we will enumerate all the possible spans within it and calculate their probabilities of being meaningful phrases, according to Equation 5. We select all the spans whose probabilities are larger than a threshold for indexing. It is worth noting that overlapping phrases are allowed at the phrase segmentation time. The selected phrases are encoded into hidden representations according to Equation 1, which will be used to construct a dense-retrieval index.
@@
For querying, we first process the query sentence (or paragraph) in a source language similarly, i.e., segmenting and encoding. Subsequently, we use the encoded query phrases to directly search for the related cross-lingual phrases using Maximum Inner Product Search (MIPS). Because of the advanced data structure and search algorithm Malkov and Yashunin (2018); Douze et al. (2024); Johnson (2022), the search step is highly efficient. The retrieved information, including the target phrase, its accompanied context, and its positions in the context, can then be integrated to downstream tasks to improve the task performance. As shown in Figure 2, the retrieved phrases along with their surrounding contexts are integrated into the instructions given to an LLM for tasks such as machine translation.
5.3 Discussion
In our preliminary studies, we also investigated various phrase segmentation strategies, including -gram segmentation and learning a Byte Pair Encoding (BPE) model Sennrich et al. (2016); Kudo (2018) across word boundaries. However, we found these methods often struggle to identify meaningful phrases, leading to significant discrepancies and performance declines between training and testing phases. Consequently, we decide to develop a trainable phrase segmentation model. In Section 6.3, we evaluate the effect of the learned phrase segmentation module on machine translation task.
6 Experiments
We propose to evaluate our model on two tasks: cross-lingual contextualized phrase retrieval and machine translation .
6.1 Implementation Details
Training Data Collection
We use all the bilingual training datasets of WMT16 on Huggingface111https://huggingface.co/datasets/wmt16 to train our Cross-lingual Contextualized Phrase Retriever (CCPR). The WMT16 dataset has six language pairs, inclduing De-En, Cs-En, Fi-En, Ru-En, and Tr-En. Because of the efficiency and the satisfactory performance of GIZA++ Och and Ney (2003); Dou and Neubig (2021), we first use the GIZA++ software222https://github.com/moses-smt/mgiza, i.e., IBM-4 model Brown et al. (1993), to induce the word alignment for each pair of parallel sentences, and then extract the cross-lingual phrase pairs. Notably, we did not induce the phrase table as in Phrase-based Statistical Machine Translation (PBSMT) Koehn et al. (2003). Instead, in our setting, we extract the cross-lingual consecutive spans that are aligned in a pair of parallel sentence as cross-lingual phrase pairs. Each of the cross-lingual phrase pairs is associated with the surrounding contexts. We filter out the phrases which begin or end with words whose frequency is more than 30k in the training dataset. We also filter out the phrases that contain only numbers and punctuations. The final training dataset contains 1.3 billion cross-lingual phrase pairs extracted from 10 million parallel sentences.
Model
In our work, we train two variants of our model, CCPR-XLMR and CCPR-LaBSE, whose parameters are initialized from XLM-RoBERTa-base (XLMR) Conneau et al. (2020) and LaBSE Feng et al. (2022), respectively. The hidden size of both models is 768. The output of in Equation (1) is 128, which plays a critical role in reducing the memory cost of our phrase index while maintaining comparable performance. Each model is trained on the mixture of all language pairs of WMT16. We train our model on 8 V100 GPUs for 20K steps, where the learning rate is 5e-5, dropout rate is 0.2, batch size on each device is 64. The for the phrase segmentation loss in Equation 5 is . Those hyper-parameters are tuned on the validation dataset. At inference time, we use the FAISS library333https://github.com/facebookresearch/faiss Douze et al. (2024) to build our retrieval index. We use the FlatIP as our index type.

### v2

@@ Abstract
Phrase-level dense retrieval has shown many appealing characteristics in downstream NLP tasks by leveraging the fine-grained information that phrases offer. In our work, we propose a new task formulation of dense retrieval, cross-lingual contextualized phrase retrieval, which aims to augment cross-lingual applications by addressing polysemy using context information. However, the lack of specific training data and models are the primary challenges to achieve our goal. As a result, we extract pairs of cross-lingual phrases using word alignment information automatically induced from parallel sentences. Subsequently, we train our Cross-lingual Contextualized Phrase Retriever (CCPR) using contrastive learning, which encourages the hidden representations of phrases with similar contexts and semantics to align closely. Comprehensive experiments on both the cross-lingual phrase retrieval task and a downstream task, i.e, machine translation, demonstrate the effectiveness of CCPR. On the phrase retrieval task, CCPR surpasses baselines by a significant margin, achieving a top-1 accuracy that is at least 13 points higher. When utilizing CCPR to augment the large-language-model-based translator, it achieves average gains of 0.7 and 1.5 in BERTScore for translations from XEn and vice versa, respectively, on WMT16 dataset. We release our code and data at https://github.com/ghrua/ccpr_release.
Cross-lingual Contextualized Phrase Retrieval
Huayang Li♣††thanks: Work done during Huayang’s internship at Tencent AI Lab. Correspondence to Deng Cai and Lemao Liu. Deng Cai Zhi Qu♣ Qu Cui Hidetaka Kamigaito♣ Lemao Liu Taro Watanabe♣ ♣Nara Institute of Science and Technology ♡Tencent AI Lab {li.huayang.lh6, qu.zhi.pv5, kamigaito.h, taro}@is.naist.jp {jcykcai, qucui, redmondliu}@tencent.com
@@ 1 Introduction
Therefore, we advocate for a new task formulation, i.e., cross-lingual contextualized phrase retrieval, which aims to find the cross-lingual phrase that is mostly relevant to the provided (general-type) source phrases, considering their meanings and surrounding contexts. However, achieving this goal is non-trivial, due to the scarcity of specific training data and models. Since annotating cross-lingual contextualized phrase pairs of general type is very difficult and expensive, we first introduce a data collection method, which leverages the automatically induced word alignment information from parallel sentences to extract suitable cross-lingual phrase pairs for training. This ensures that the phrases are of general type and both cross-lingual phrases and contexts are well aligned. Thereafter, we propose a Cross-lingual Contextualized Phrase Retriever (CCPR) which is trained on the constructed dataset. The CCPR mainly employs phrase-level contrastive learning to draw cross-lingual phrases with similar contexts and meanings closer in the hidden space. At inference time, we can leverage the well-trained CCPR to build a phrase-level index and use the query phases for search.
Unlike previous works Zheng et al. (2022), which focus solely on the task of cross-lingual phrase retrieval for wiki entities, our work also explores the potential of leveraging CCPR to augment downstream cross-lingual tasks, e.g., machine translation (MT). One critical question in front of us is how to select meaningful phrases for indexing and searching at inference time. To address this problem, we propose to learn a phrase segmentation module, which can be used to predict meaningful phrases from sentence- or passage-level retrieval data at inference time. The developed phrase segmentation module is important for ensuring the train-test consistency of CCPR.
@@
Experiments on both of the cross-lingual contextualized phrase retrieval and a downstream task, i.e., MT, show the effectiveness of CCPR. We first evaluate CCPR on the cross-lingual contextualized phrase retrieval task, since a higher accuracy is more beneficial for downstream tasks, such as MT, cross-lingual dictionary induction, etc Zhang and Zong (2016); Søgaard et al. (2018). For this task, we use the human annotated cross-lingual phrase pairs as the golden truth and evaluate whether CCPR and other baselines can retrieve those golden phrases from a large-scale index. Experiments show that CCPR outperforms the baselines for at least 13 points in terms of the top-1 accuracy. We also conduct evaluation on the MT task using a large language model (LLM), e.g., LLama-2 Touvron et al. (2023). This task aims to evaluate whether the information retrieved by CCPR can enhance LLM’s ability of cross-lingual generation. Following the fashion of retrieval-augmented generation (RAG) Lewis et al. (2020b), we simply integrate the retrieved phrase information to the input of the LLM, and compare our method with other baselines. For both XEn and EnX translation directions on the WMT16 dataset, where X is from six languages, our CCPR achieves averaged gians of 0.7 and 1.5 BERTScore points Zhang et al. (2019), respectively. Results on other evaluation metrics, e.g., COMET Rei et al. (2022), are consistent with BERTScore.
@@
We propose a Cross-lingual Contextualized Phrase Retriever (CCPR), which uses the constructed training data, i.e., cross-lingual phrases extracted from automatically induced word alignment information, to learn both the phrase alignment and segmentation modules.
We conduct extensive experiments on cross-lingual contextualized phrase retrieval and MT. Our method outperforms the baselines by a large margin.
@@ 2 Related Work
In text generation, dense retrieval has been widely studied at both sentence and phrase levels. Particularly after the emergence of large language models (LLMs), how to retrieve related sentences (or passages) Izacard et al. (2021); Ni et al. (2022) for input to LLMs has became a popular research topic, namely, retrieval-augmented generation (RAG) Lewis et al. (2020b); Karpukhin et al. (2020); Guu et al. (2020); Borgeaud et al. (2022); Asai et al. (2024). However, compared to the sentence-level retrieval, phrase-level retrieval has been shown to not only enhance retrieval accuracy but also markedly boost performance in downstream tasks Lee et al. (2021a, b). By indexing and integrating retrieved phrases into the output of language models, Lan et al. (2023) and Cao et al. (2024) make the text generation process more attributable and accurate. These studies underline the effectiveness of phrase-level dense retrieval. Our work notably diverges from those works, because we focus more on cross-lingual tasks.
@@
In cross-lingual tasks, the majority of research has concentrated on sentence-level retrieval. Many research works in retrieval-augmented MT Zhang et al. (2018); Gu et al. (2018); Xia et al. (2019); He et al. (2021) have explored retrieving bilingual sentences based on similarity of source sentences, guiding MT models to utilize retrieved target information. A notable limitation is their reliance on bilingual data for constructing the index. This requirement substantially limits the scale and diversity of data available for retrieval. In contrast, Cai et al. (2021) propose to directly retrieve relevant target sentences using source sentences to assist MT, facilitating cross-lingual retrieval using a monolingual index. Unlike the task-specific model for MT in Cai et al. (2021), many studies in dense retrieval aim to enhance the cross-lingual sentence retrieval on more languages and at a larger scale Chidambaram et al. (2019); Conneau et al. (2020); Reimers and Gurevych (2020); Heffernan et al. (2022); Feng et al. (2022); Cai et al. (2022).
@@
Additionally, there is research focused on employing nearest neighbor retrieval to support MT Khandelwal et al. (2021); Zheng et al. (2021); Meng et al. (2022); Deguchi et al. (2023). These efforts differ from the aforementioned studies as they aim to estimate prediction distributions through nearest neighbor retrieval at each translation step.
@@ 3 Task Formulation
This work aims a cross-lingual phrase retriever that can augment downstream NLP tasks. To ensure its effectiveness, the retriever must adhere to two essential criteria. First, it should ensure that phrases in the index are of general type. Second, it must resolve the polysemy of general-type phrases by considering their contextual differences.
@@
This new task formulation also introduces several challenges, with the primary issue being the lack of training data. Specifically, there is a scarcity of cross-lingual phrase pairs of general type and accompanied by contextual information. Such data are crucial for training the model to recognize cross-lingual phrases that have similar contexts and meanings, allowing it to align these phrases closely in a hidden space. In addition, how to extract meaningful general-type phrases from sentences is also an open question, which is critical for employing the retriever to downstream tasks.
@@ 4 Training Data Collection
Formally, given a pair of parallel sentences , and , where and are the number of words in the sequences, we use a word alignment model to obtain the word alignment information of and Koehn et al. (2003); Och et al. (1999). If every word in a consecutive span can be aligned to a consecutive span , we will use them as a pair of aligned phrases and , where and . Notably, phrases that meet the constraints may range from a single word to the whole sequence, and extracted phrases are allowed to have overlaps. In addition, the context information of and are and , respectively. Unlike the phrase table extracted in SMT Koehn et al. (2003); Chiang (2005), which is static and context-independent, here the phrase mapping between and is unique considering their context.
@@ 5.1 Model Architecture
We introduce a Cross-lingual Contextualized Phrase Retriever (CCPR) for our new formulated task. The main target of CCPR is to make the representations of cross-lingual phrases with similar contexts and meanings to be close in the hidden space. To this end, we propose a cross-lingual phrase alignment module based on contrastive learning, utilizing the data collected in Sec. 4 for training. In addition, one remaining problem is how to select phrases for indexing at inference time. To address this, we propose a phrase segmentation module that uses phrase representations to predict meaningful phrases from sentences or paragraphs. At inference time, we can leverage the learned phrase segmentation module to select phrases for indexing, ensuring the train-test consistency.
@@
where is a Transformer model, e.g., BERT Devlin et al. (2019); Feng et al. (2022) and RoBERTa Conneau et al. (2020), that encodes the context to a matrix with a dropout mask . Inspired by preview works Lan et al. (2023); Lee et al. (2021a); Seo et al. (2018), we use the concatenation of and , i.e., the start and end token of the phrase, to represent phrase . is a function that maps the concatenation of two hidden states and from to , where is the output hidden size. The encoding for phrases in context is similar.
@@
For a batch of parallel sentences, we can extract a collection of cross-lingual phrase pairs , where and are the positive examples of each other. Since each positive phrase pair is from parallel sentences that share similar semantics, directly learning on them may cause the model to learn a trivial shortcut. Therefore, inspired by SimCSE Gao et al. (2021), we apply two independently sampled dropout mask and , to encode phrases and , respectively, following Eq. 1. The training objective for the direction is:
where we use all the non-paired phrases in as in-batch negatives in Eq. 3. For bidirectional symmetry, our final loss is:
The illustration of the contrastive learning is shown in Fig. 1 (a). In practice, we may use parallel data from multiple languages during training, i.e., the source contexts in a batch may come from multiple languages, e.g., Germany, Czech, etc, which is inspired by the success of multi-lingual training in Liu et al. (2020).
@@
This module aims to learn how to select phrases from sentences or passages, since most retrieval data at inference time comes in sentences or passages, rather than phrases. To ensure the train-test consistency, we take the phrases extracted in Sec. 4 as positive data, and all the other spans in the corresponding contexts as negative data. Our phrase segmentation module is a binary classifier defined as follows:
where is the label for the span , where , is an activation function, is linear layer that maps the hidden representation from to , and the definitions for and are the same as in Eq. 1. We use the label for phrases and for non-phrase spans. The phrase segmentation for sequence is similar to Eq. 5. In practice, the number of non-phrase spans is significantly more than the number of phrases. To mitigate data imbalance, we employ a strategy of randomly sampling an equal number of non-phrase spans and phrases within a sentence during training Li et al. (2020). The illustration of phrase segmentation is in Fig. 1 (b).
@@
In our preliminary studies, we investigated various phrase segmentation strategies, including -gram segmentation and learning a Byte Pair Encoding (BPE) model Sennrich et al. (2016); Kudo (2018) across word boundaries. However, these methods struggle to identify meaningful phrases, leading to significant discrepancies between training and testing phases and performance declines. In Sec. 6.3, we evaluate the effect of the learned phrase segmentation on MT task.
@@ 5.2 Inference Pipeline
The pipeline of employing our cross-lingual contextualized phrase retriever for downstream NLP tasks involves two main steps: index building and searching, as shown in Fig. 1 (c).
@@
A significant benefit of our retriever is its ability to leverage monolingual data in the target language, eliminating the need for bilingual data to construct the index at inference time Zhang et al. (2018); Gu et al. (2018); Bapna and Firat (2019). Given the retrieval data in target language, our retriever begins by segmenting sentences (or passages) into phrases. When segmenting a sentence (or passage), we will enumerate all the possible spans within it and calculate their probabilities of being meaningful phrases, according to Eq. 5. We select all the spans whose probabilities are larger than a threshold for indexing. It is worth noting that overlapping phrases are allowed at the phrase segmentation time. The selected phrases are encoded into hidden representations according to Eq. 1, which will be used to construct a dense-retrieval index.
@@
For querying, we first process the query sentence (or paragraph) in a source language similarly, i.e., segmenting and encoding. Subsequently, we use the encoded query phrases to directly search for the related cross-lingual phrases using Maximum Inner Product Search (MIPS). Because of the advanced data structure and search algorithm Malkov and Yashunin (2018); Douze et al. (2024); Johnson (2022), the search step is highly efficient (more details in App. A.3). The retrieved information, including the target phrase, its accompanied context, and its positions in the context, can then be integrated to downstream tasks to improve the performance. As shown in Fig. 2, the retrieved phrases along with their surrounding contexts are integrated into the instructions given to an LLM for tasks such as MT.
@@
6 Experiments
We propose to evaluate our model on cross-lingual contextualized phrase retrieval and MT.
6.1 Implementation Details
Training Data Collection
We use all the bilingual training datasets of WMT16 on Huggingface111https://huggingface.co/datasets/wmt16 to train our Cross-lingual Contextualized Phrase Retriever (CCPR). The WMT16 dataset has six language pairs, inclduing De-En, Cs-En, Fi-En, Ru-En, and Tr-En. Because of the efficiency and the satisfactory performance of GIZA++ Och and Ney (2003); Dou and Neubig (2021), we first use the GIZA++ software222https://github.com/moses-smt/mgiza, i.e., IBM-4 model Brown et al. (1993), to induce the word alignment for each pair of parallel sentences, and then extract the cross-lingual phrase pairs. More details about phrase extraction are in App. A.1. The final training dataset contains 1.3 billion cross-lingual phrase pairs extracted from 10 million parallel sentences.
Model
In our work, we train two variants of our model, CCPR-XLMR and CCPR-LaBSE, whose parameters are initialized from XLM-RoBERTa-base (XLMR) Conneau et al. (2020) and LaBSE Feng et al. (2022), respectively. The hidden size of both models is 768. The output of in Eq. (1) is 128, which plays a critical role in reducing the memory cost of our phrase index while maintaining comparable performance. At inference time, we use the FAISS library333https://github.com/facebookresearch/faiss Douze et al. (2024) to build our retrieval index. We use the FlatIP as our index type. Training details are App. in A.2